OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 8x - (5x + 4) greater than or equal to -31?

OpenStudy (anonymous):

need help please.

OpenStudy (anonymous):

Where's the answer choices?

OpenStudy (anonymous):

x greater than or equal to -9 -9 less than or greater to x A number line with a closed circle on -9 and shading to the left A number line with a closed circle on -9 and shading to the right

OpenStudy (texaschic101):

8x - (5x + 4) >= -31 8x - 5x - 4 >= - 31 3x >= -31 + 4 3x >= - 27 x >= - 9 or -9 < = x number line with a closed circle on -9 (because if it has an equal sign, it has a closed circle) and it is shaded to the right (because it is greater then). so what do you think the answer is ?

OpenStudy (anonymous):

d

OpenStudy (texaschic101):

no...try again

OpenStudy (anonymous):

c

OpenStudy (texaschic101):

yep..that one is not true. Because the shading will go to the right, not the left

OpenStudy (anonymous):

oh i see i have one more it similar to it if you don't mind.

OpenStudy (texaschic101):

I don't mind

OpenStudy (anonymous):

Choose the correct description of the graph of the compound inequality 3x + 2 > 2 and 3x less than or greater to 6.

OpenStudy (texaschic101):

is that supposed to be less then or equal ?

OpenStudy (anonymous):

then the answer: A number line with an open circle on 0, shading to the left, and a closed circle on 2, shading to the right A number line with a closed circle on 0, shading to the left, and an open circle on 2, shading to the right A number line with a closed circle on 0, an open circle on 2, and shading in between A number line with an open circle on 0, a closed circle on 2, and shading in between

OpenStudy (anonymous):

less then

OpenStudy (texaschic101):

3x + 2 > 2 3x > 2 - 2 3x > 0 x > 0 3x <= 6 x <= 2 open circle on 0, closed circle on 2, shading in between

OpenStudy (texaschic101):

remember this.. open circle if there is no equal sign closed circle if there is an equal sign < less then means shading to the left > greater then means shading to the right

OpenStudy (anonymous):

so it would be D

OpenStudy (texaschic101):

yep...its D